# ConCISE: Confidence-guided Compression in Step-by-step Efficient Reasoning

### v1

@@ 4.2.4 Can reasoning compression methods generalize across tasks?
Math-500 Origin 90.8 3854.2 113.7 33.9 100%
ConCISE 92.0 2244.3 63.0 35.6 58.2%
Conf. Inj. 91.8 2619.6 79.5 33.0 68.0%
Early Stop 92.4 2653.8 68.4 38.8 68.9%
GSM8K Origin 93.1 1441.9 41.4 34.8 100%
ConCISE 92.9 831.9 22.5 37.1 57.7%
Conf. Inj. 92.6 933.8 23.2 40.2 64.8%
Early Stop 92.7 1003.0 22.3 45.0 69.6%
AIME24 Origin 54.2 13573.5 463.0 29.3 100%
ConCISE 52.1 9750.8 333.0 29.3 71.8%
Conf. Inj. 51.3 10165.8 372.2 27.3 74.5%
Early Stop 54.2 12204.6 389.7 31.3 89.9%
Table 2: Ablation study of DeepSeek-R1-Distill-Qwen-7B across three mathematical benchmarks.
4.3 Ablation Study
We conduct an ablation study to evaluate the two core components of ConCISE: Confidence Injection and Early Stopping. For each mechanism, we construct a training dataset and fine-tune the base model using the same configurations as in the main experiments. All experiments are performed in the supervised fine-tuning (SFT) setting and evaluated across three mathematical reasoning tasks: Math-500, GSM8K, and AIME24. The results, summarized in Table 2, show that both mechanisms achieve similar reasoning accuracy to full ConCISE but exhibit weaker compression performance. Both methods focus on reducing redundant reflection steps while maintaining the integrity of the reasoning process.
Confidence Injection suppresses unnecessary iterations by reinforcing the model’s confidence, but it has a limited impact on terminating reasoning after the final answer. Early Stopping, on the other hand, halts excessive post-answer reasoning but does not address earlier stages. Interestingly, Early Stopping nearly preserves the original model’s accuracy on benchmarks like AIME24, suggesting that post-FAS reflection steps often contain substantial redundancy and can be safely stopped without harming task performance. In summary, Confidence Injection reduces redundancy across the entire reasoning sequence, while Early Stopping targets post-answer overthinking. Only by combining both can ConCISE achieve comprehensive redundancy elimination, resulting in superior compression without compromising reasoning quality.
@@ 5 Conclusion
In this work, we first modeled the generation of reflection steps in Large Reasoning Models (LRMs) from a confidence perspective. This modeling approach enabled us to systematically identify two primary patterns of redundancy: Confidence Deficit, where models unnecessarily undertrust intermediate steps, and Termination Delay, where reflection persists despite reaching a confident answer. To address these issues, we introduced ConCISE, a novel confidence-guided framework designed to enhance reasoning efficiency. ConCISE employs two targeted mechanisms: Confidence Injection to suppress redundant reflections by boosting internal confidence, and Early Stopping to halt generation upon detecting sufficient answer confidence. Our experiments empirically validate that ConCISE significantly compresses response length (~50% under SimPO) while preserving strong accuracy across diverse reasoning benchmarks. Notably, ConCISE achieves superior compression-accuracy trade-offs over existing methods and exhibits robust generalization to out-of-domain tasks. This study offers a principled approach to modeling and mitigating redundancy in LRM reasoning, enabling more resource-efficient and user-friendly AI systems.
@@
Further Compression of Step Length.
While ConCISE effectively targets redundant reflection steps, analysis indicates substantial potential remains for intra-step compression based on average step length. Future work could explore integrating techniques like TokenSkip with ConCISE for greater compression.
@@
Although ConCISE introduces a lightweight confidence detector for the post-answer phase, it currently lacks a direct mechanism to model confidence prior to the first answer, relying instead on reflection occurrences as an indirect proxy.
Model Dependency and Sensitivity.
Validated on DeepSeek models, ConCISE’s effectiveness and the optimal tuning of its components (e.g., confidence phrases, detector, thresholds) may be sensitive to model architecture or scale. Its performance across diverse LRMs requires further study, potentially necessitating model-specific adaptations.
@@ A.2 Early Stopping
probing prompt and confidence-indicated tokens. The effectiveness of the Early Stopping mechanism relies heavily on the accuracy of its confidence detector. To identify an appropriate probing prompt, we analyzed the reasoning chains generated by the DeepSeek-R1-Distill-Qwen-7B model using Greedy decoding on the training question set. This analysis revealed that the model frequently expresses confidence during the later part of the reasoning chain. Table 4 summarizes the frequencies of the most common self-reported confidence expressions observed within chains.
Table 4: Frequency of Common Confidence Expressions observed in 2000 Reasoning Chains.
Confidence Expression Frequency Count
I’m confident 423
I feel confident 170
I can be confident that 157
I can confidently say 70
This gives me confidence that 68
I’m pretty confident 43
I’m pretty sure 37
Based on the prevalence of phrases starting with "I’m" leading into confidence statements, we selected "So, I’m" as the fixed probing prompt. Subsequently, we identified "confident", "pretty confident", "sure", and "pretty sure" as the primary confidence-indicative tokens.
@@
[…]
@@
Early stopping threshold. The selection of the early stopping threshold, $t$, was guided by the principle of tailoring the reflection process to the apparent quality of the reasoning chain. Our objective was to configure the threshold such that:
@@
To find a threshold value that aligns with this principle, we empirically evaluated $t_{e}\in\{0.4,0.5,0.6,0.7\}$. For each value, we measured the probability distribution of the stopping points across different stages: stopping after the initial answer generation (0 reflections), after the first reflection, after the second reflection, etc. The results of this analysis are illustrated in Figure 6.
Figure 6: Distribution of early stopping points (probability of stopping at Pre-FAS, FAS, Ref1, 2, …, 5) for different threshold values ($t_{e}$).
Observing the distributions in Figure 6, we found that setting the threshold $t_{e}=0.5$ yielded the stopping behavior that best matched our desired criteria across varying problem difficulties. Therefore, we selected $t_{e}=0.5$ as the definitive early stopping threshold for all subsequent experiments conducted in this work.
@@ A.3 Reflection step detection
[…]
@@
  • "wait", "alternatively", "check", "reconsider", "reflect", "rethink", "reconsidering", "reviewing", "reassess", "pause", "second thought", "reevaluate", "verify".
@@ A.3.2 LLM-as-a-Judge
Furthermore, the LLM-as-a-Judge approach is the basis for calculating several reflection-based metrics presented in our results, such as RefNum, Non-RefNum, pre-FAS RefNum, and post-FAS RefNum (shown in Figure 5(a) and Figure 5(b)).
@@ Appendix B Details about Experiments
B.1 Question Set
Figure 7: Difficulty diversity of question set.
Figure 8: Length diversity of question set.
While the construction methodology for the question set is detailed in Section 4, here we focus on illustrating its diversity in terms of difficulty and length.Figures 7 and 8 demonstrate the diversity of the Question set. Difficulty distribution is derived from metric Level in the MATH dataset, while length distribution reflects reasoning chain lengths from DeepSeek-R1-Distill-Qwen-7B under greedy decoding.
B.2 Validation Dataset
@@
B.3 Evaluation Dataset
The analysis presented in Figure 5 required LLM-as-a-Judge annotations. To ensure reliable results despite this method’s limitations with long contexts, we filtered the MATH500 dataset. We selected only those problems where reasoning chains generated by all twelve specified finetuning configurations (combinations of ConCISE/OverThink/Spirit methods, 1.5B/7B sizes, and SFT/SimPO techniques) were shorter than 5000 tokens. This yielded a common subset of 340 problems, on which our subsequent comparative analysis of the different methods was exclusively performed.

### v2

@@ 4.3 Ablation Study
Math-500 Origin 90.8 3854 113.7 33.9 100%
ConCISE 92.0 2244.3 63.0 35.6 58%
Conf. Inj. 91.8 2620 79.5 33.0 68%
Early Stop 92.4 2654 68.4 38.8 69%
GSM8K Origin 93.1 1442 41.4 34.8 100%
ConCISE 92.9 832 22.5 37.1 58%
Conf. Inj. 92.6 934 23.2 40.2 65%
Early Stop 92.7 1003 22.3 45.0 70%
AIME24 Origin 54.2 13574 463.0 29.3 100%
ConCISE 52.1 9751 333.0 29.3 72%
Conf. Inj. 51.3 10166 372.2 27.3 75%
Early Stop 54.2 12205 389.7 31.3 90%
Table 2: Ablation study of Confidence Injection and Early Stopping on DeepSeek-R1-Distill-Qwen-7B.
We conduct an ablation study to evaluate the two core components of ConCISE: Confidence Injection and Early Stopping. For each, we construct a training dataset and fine-tune DeepSeek-7B using the same configurations as in the main experiments. All experiments are performed under SFT settings and evaluated across three mathematical reasoning tasks: Math-500, GSM8K, and AIME24.
The results, summarized in Table 2, show that both mechanisms achieve similar reasoning accuracy to full ConCISE but exhibit weaker compression performance. Specifically, Confidence Injection suppresses unnecessary reflections by boosting the model’s confidence, but has a limited impact on terminating reasoning after the final answer. Early Stopping, on the other hand, halts excessive post-answer reasoning but does not address earlier stages. Only by combining both can we achieve comprehensive redundancy elimination without compromising reasoning quality.
Figure 6: Average performance of training-free methods and ConCISE variants on DeepSeek-R1-Distill-Qwen-7B and 1.5B across GSM8K, Math-500, and AIME24.
4.4 Analysis of Training-free Methods
To further demonstrate the advantages of the ConCISE framework, we compare it against representative training-free methods. Specifically, we include TALE Han et al. (2024), which constrains reasoning length by injecting explicit token budgets into prompts, and DEER Yang et al. (2025), which dynamically decides when to terminate reasoning by monitoring the generation probability of the answer. For DEER, we reproduce two variants with $\lambda=0.95$ and $\lambda=0.98$. We also implement a training-free version of our method, denoted ConCISE-Decoding, which directly applies Algorithm 1during inference to decide when to terminate reasoning, without relying on ground-truth verification. We evaluate these methods on GSM8K, Math-500, and AIME24, and report the averaged results in Figure 6, with detailed statistics provided in Table 6.
We find that training-free methods can achieve partial compression, but their compression levels are unstable and often come with substantial accuracy degradation. For instance, DEER achieves relatively stable compression but introduces a 5–8% accuracy drop. Meanwhile, TALE attains a high compression ratio on GSM8K with DeepSeek-7B, but performs poorly on other datasets and models. In contrast, ConCISE-SimPO consistently achieves a superior overall trade-off, and even the training-free variant ConCISE-Decoding outperforms other training-free approaches. These results highlight that ConCISE enables more precise elimination of redundant reflection steps, thereby yielding higher-quality reasoning chains.
@@ 5 Conclusion
In this work, we first modeled the generation of reflection steps in LRMs from a confidence-guided perspective, helping to understand two key patterns responsible for redundant reflection within the reasoning process of LRMs: Confidence Deficit and Termination Delay. Based on this, we propose ConCISE, which employs Confidence Injection and Early Stopping mechanisms to respectively address the above two patterns to suppress redundant reflection steps and produce efficient, concise reasoning chains. Extensive experiments demonstrate that LRMs fine-tuned on ConCISE-generated data significantly compress their responses while maintaining strong task performance.
@@
Room for Further Compression.
While ConCISE effectively addresses redundant reflection steps, analysis indicates that significant potential remains for further compression of both non-reflection steps and the average length of each step. Future work could explore integrating techniques with ConCISE to achieve a higher degree of compression while maintaining model performance.
@@
Although ConCISE introduces a lightweight confidence detector for the post-answer phase, it still lacks a direct mechanism to model confidence before the first answer. Instead, it relies on the occurrence of reflection steps as an indirect proxy. Future work could explore training a lightweight model to detect the model’s internal confidence, thereby enabling more fine-grained control.
Integration with RLVR methods.
This work focuses on constructing concise reasoning traces and validating them under SFT and SimPO, which we consider sufficient to support our core contribution. At the same time, we recognize that RLVR methods such as GRPOShao et al. (2024) are powerful and widely used in post-training. Exploring how confidence-guided signals can be incorporated into RLVR setups is a valuable future direction.
Acknowledgement
This research was supported in part by the National Natural Science Foundation of China under Grant No. 62402267, 62432004, the China Postdoctoral Science Foundation under Grant No. BX20240177, 2025M771499, and a grant from the Guoqiang Institute, Tsinghua University.
@@ A.1 Confidence Injection
Additionally, another crucial mechanism within the Confidence Injection process is identifying whether the current step constitutes a reflection step. The accurate and swift identification of such steps is paramount for the success of Confidence Injection. Consequently, informed by our observations of reflection steps during LRM reasoning processes, we employed a rule-based detection method. This method, which is detailed in Appendix A.3.1, achieves reasonably accurate detection while having a negligible impact on the model’s inference speed.
@@ A.2 Early Stopping
Probing prompt and confidence-indicated tokens.
The effectiveness of the Early Stopping mechanism relies heavily on the accuracy of its confidence detector. To identify an appropriate probing prompt, we first analyzed the reasoning chains generated by four Large Reasoning Models (LRMs) using Greedy decoding on the training question set. This analysis revealed that the model frequently expresses confidence during the later part of the reasoning chain, and as shown in Table 4, self-reported confidence expressions starting with “I’m” (such as “I’m confident”) are very common.
Based on this observation, particularly considering the prevalence of confidence statements led by “I’m”, we considered "So, I’m" as a potential probing prompt. Concurrently, we also experimented with alternative probing prompts, such as "So, I feel" and "I can be". However, we observed that these alternative prompts often resulted in high confidence scores before the model had reached the final answer (i.e., prior to Final Answer Serialization, FAS), leading to unnecessary computational overhead.
Therefore, considering these factors—namely, the prevalence of “I’m”-led confidence statements in the model’s natural expressions and the shortcomings of the alternative prompts—we ultimately selected "So, I’m" as the fixed probing prompt. Subsequently, we identified the primary confidence-indicative tokens that follow the "So, I’m" prompt, which are: "confident", "pretty confident", "sure", and "pretty sure".
Table 4: Frequency of Common Confidence Expressions Observed in 2000 training Reasoning Chains Across Different Large Reasoning Model Series(DeepSeek, Qwen and Skywork)
Expression DeepSeek-7B DeepSeek-1.5B Qwen3-8B Skywork-OR17B
I’m confident 423 335 127 560
I feel confident 170 172 58 144
I can be confident that 157 156 475 130
I can confidently say 70 176 91 98
This gives me (more)confidence that 68 81 23 86
I’m pretty confident 43 69 2 60
I’m pretty sure 37 28 2 17
Therefore, confident that 1 0 26 5
Total (sum of listed expressions) 969 1017 804 1100
@@
$\hat{c}_{i}=$ $P(\text{"confident"}\mid S_{i},\text{"So, I'm"})$
$+P(\text{"sure"}\mid S_{i},\text{"So, I'm"})$
$+P(\text{"pretty"}\mid S_{i},\text{"So, I'm"})\times$
$\big(P(\text{"confident"}\mid S_{i},\text{"So, I'm pretty"})$
$\quad+P(\text{"sure"}\mid S_{i},\text{"So, I'm pretty"})\big)$
@@
Early stopping threshold.
The selection of the early stopping threshold, $t$, was guided by the principle of tailoring the reflection process to the apparent quality of the reasoning chain. Our objective was to configure the threshold such that:
@@
To find a threshold value that aligns with this principle, we empirically evaluated $t_{e}\in\{0.4,0.5,0.6,0.7\}$ on the DeepSeek-R1-Distill-Qwen-7B model. For each value, we measured the probability distribution of the exceeding points(when $\hat{c}_{i}>t_{e}$) across different stages: exceeding after the initial answer generation (0 reflections), after the first reflection, after the second reflection, etc. The results of this analysis are illustrated in Figure 7.
Figure 7: Distribution of exceeding points(when $\hat{c}_{i}>t_{e}$) probability of Pre-FAS, FAS, Ref1, 2, …, 5 for different threshold values ($t_{e}$).
Benchmark $T_{e}$ Acc. Tok. StepNum StepLen CR
Math-500 0.4 91.4 2390.6 68.4 34.9 62.0%
0.5 92.0 2244.3 63.0 35.6 58.2%
0.6 91.8 2324.3 66.1 35.2 60.3%
0.7 92.0 2345.1 67.5 34.8 60.8%
GSM8K 0.4 92.4 828.9 19.9 41.6 57.5%
0.5 92.9 831.9 22.5 37.1 57.7%
0.6 93.0 831.7 20.4 40.7 57.7%
0.7 92.7 849.3 21.7 39.2 58.9%
AIME24 0.4 45.8 11548.7 402.1 28.7 85.1%
0.5 52.1 9750.8 333.0 29.3 71.8%
0.6 52.1 9802.5 340.5 28.8 72.2%
0.7 51.7 9719.9 339.8 28.6 71.6%
Table 5: Effect of different early stopping thresholds (0.4, 0.5, 0.6, 0.7) on the SFT performance of DeepSeek-R1-Distill-Qwen-7B across mathematical benchmarks.
Observing the distributions in Figure 7, we found that these $t_{e}$ settings generally cause the model to achieve a confidence score $\hat{c}_{i}>t_{e}$ at the position of the first reflection, which aligns with our three guiding principles. Furthermore, the confidence monitored by our designed lightweight detector rarely indicates high confidence before the FAS (Final Answer Serialization) step. This is advantageous as it effectively reduces the computational overhead associated with prematurely judging the correctness of an answer, making it highly compatible with the ConCISE methodology.
Therefore, we further conducted experiments for $t_{e}\in\{0.4,0.5,0.6,0.7\}$ under the SFT setting of DeepSeek-R1-Distill-Qwen-7B, and Table 5 presents these experimental results. We found that under the settings of $t_{e}=0.5,0.6,$ and $0.7$, the model ultimately exhibited similar performance in both compression rate (CR) and accuracy (Acc), demonstrating robustness to the choice of $t_{e}$. Consequently, we selected $t_{e}=0.5$ for subsequent extensive experiments. However, when $t_{e}=0.4$, there was a significant decline in performance. This is because when $t_{e}$ is set to $0.4$, more data exists at the FAS step, preventing the model from adequately checking its reasoning process; this excessive confidence leads to a performance drop.
@@ A.3 Reflection step detection
Formally, let $K$ be the predefined set of reflection keywords. A step $s_{i}$ is identified as the start of a reflection, denoted as $\text{IsReflectionStart}(s_{i})$, if its text contains any keyword $k$ from the set $K$:
$\text{IsReflectionStart}(s_{i})\equiv\bigvee_{k\in K}(k\subseteq s_{i})$
@@
  • "wait", "alternatively", "check", "reconsider", "reflect", "rethink", "reconsidering", "reviewing", "reassess", "pause", "second thought", "reevaluate", "verify", "think again".
@@
(a) Difficulty diversity of question set.
(b) Length diversity of question set.
Figure 8: Diversity of the question set, showing distributions for difficulty and length.
@@ A.3.2 LLM-as-a-Judge
Furthermore, the LLM-as-a-Judge approach is the basis for calculating several reflection-based metrics presented in our results, such as RefNum, Non-RefNum, pre-FAS RefNum, and post-FAS RefNum (shown in Figure 5(b) and Figure 5(c)).
Model Method GSM8K Math-500 AIME24 Average
Acc. CR Acc. CR Acc. CR Acc. CR
DeepSeek-7B Origin 93.1 100% 90.8 100% 54.2 100% 79.4 100%
DEER(0.98) 90.3 59% 87.0 51% 47.1 69% 74.8 60%
DEER(0.95) 87.9 56% 85.2 49% 43.4 64% 72.2 56%
TALE 89.0 47% 90.0 95% 52.9 98% 77.3 80%
ConCISE-Decoding 92.7 62% 89.2 55% 49.6 67% 77.2 61%
ConCISE-SimPO 92.1 50% 91.0 51% 48.3 57% 77.1 53%
DeepSeek-1.5B Origin 85.4 100% 82.2 100% 29.2 100% 65.6 100%
DEER(0.98) 79.7 41% 77.6 54% 24.6 70% 60.6 55%
DEER(0.95) 74.7 43% 76.4 49% 21.3 64% 57.5 52%
TALE 84.9 90% 81.6 98% 29.6 98% 65.4 95%
ConCISE-Decoding 83.9 53% 79.0 52% 26.3 52% 63.1 52%
ConCISE-SimPO 84.3 36% 83.6 51% 30.4 50% 66.1 46%
Table 6: Detailed results of training-free baselines and ConCISE variants on DeepSeek-R1-Distill-Qwen-7B and 1.5B across GSM8K, Math-500, and AIME24. Acc: accuracy (%), CR: compression ratio relative to original (lower CR indicates shorter outputs).
(a) Average Acc, StepNum, and StepLen.
(b) Analysis of Reflection steps.
(c) Metrics of training datasets.
Figure 9: Further analysis of reasoning chain and training dataset characteristics on DeepSeek-R1-Distill-Qwen-1.5B.
@@ Appendix B Details about Experiments
B.1 Datasets
Question Set
While the construction methodology for the question set is detailed in Section 4, here we focus on illustrating its diversity in terms of difficulty and length.Figures 8(a) and 8(b) demonstrate the diversity of the Question set. Difficulty distribution is derived from metric Level in the MATH dataset, while length distribution reflects reasoning chain lengths from DeepSeek-R1-Distill-Qwen-7B under greedy decoding. Furthermore, it is important to note a subsequent refinement process for these selected questions when constructing the final training set. Although the initial 2000 data entries were chosen based on the premise that the model could correctly answer them using greedy search, the generation processes for our ConCISE and OverThink methods do not strictly adhere to greedy decoding, and ConCISE will discard wrong reasoning chains. Therefore, for training purposes, we ultimately utilize the common subset of correctly reasoned chains derived from the outputs of three methods: ConCISE, OverThink, and Spirit. This intersection results in a refined training dataset of approximately 1900 samples.
Validation Dataset
@@ B.1 Datasets
Evaluation Dataset
The analysis presented in Figure 5(b) and Figure 9(b) required LLM-as-a-Judge annotations. To ensure reliable results despite this method’s limitations with long contexts, we filtered the MATH500 dataset. We selected only those problems where reasoning chains generated by all twelve specified finetuning configurations (combinations of ConCISE/OverThink/Spirit methods, 1.5B/7B sizes, and SFT/SimPO techniques) were shorter than 5000 tokens. This yielded a common subset of 340 problems, on which our subsequent comparative analysis of the different methods was exclusively performed.
B.2 Training Details
Models.
DeepSeek-R1-Distill-Qwen-7B and DeepSeek-R1-Distill-Qwen-1.5B are reasoning models derived by DeepSeek from the Qwen2.5 series models Yang et al. (2024)through distillation training, using training and generation data from the DeepSeek-R1 model. Skywork-OR1-7B-Preview was further developed via reinforcement learning, exhibiting strong capabilities on test sets such as mathematics. Meanwhile, Qwen3-8B is a recently released hybrid reasoning model noted for its powerful reasoning abilities. All four of these are open-source models. While they possess strong Chain-of-Thought (CoT) reasoning capabilities, they also present clear issues with reasoning redundancy. Consequently, we employ these four models for subsequent training and evaluation. Furthermore, we ensure that these models, along with all datasets involved in the training and evaluation processes, are utilized following their original licenses and intended purposes.
Training.
We construct training datasets based on the question set using ConCISE, OverThink, and Spirit. For the generation process of ConCISE, we set Temperature=0.6, Top_p=0.95. For OverThink, we sample 8 reasoning chains under Tempature=1.0(which is relatively high due to the recommended temperature for solving math problems of the LRMs is usually 0.6-0.8), Top_p=0.95. For Spirit, compression is applied to reasoning chains generated via greedy search. During training, we adopt two strategies: SFT and SimPO. SFT directly uses the question and its corresponding compressed reasoning trace for supervised fine-tuning. For SimPO, reject samples are also required Rafailov et al. (2023). Following the original OverThink setup, we use the longest correct reasoning chain among the eight samples as the reject samples. To ensure a fair comparison, ConCISE adopts the same strategy. For Spirit, the full reasoning chain before compression serves as reject samples, because this is precisely the comparative effect Spirit aims to demonstrate: important steps are retained while preference relationships are learned only for non-important steps. We use LlamaFactory for both SFT and SimPO fine-tuning Zheng et al. (2024). All models are trained with lr=1e6lr=1e-6; SFT runs for 2 epochs and SimPO for 1 epoch. The effective batch size for all training is 32 (using per_device_train_batch_size=1 and gradient_accumulation_steps=8 on 4 A800 GPUs), and max_length is set to 10240.
B.3 Analysis on DeepSeek-1.5B
Regarding the analysis conducted on DeepSeek-R1-Distill-Qwen-7B in the Experimental Section 4.1, we also performed a similar analysis for DeepSeek-R1-Distill-Qwen-1.5B. Figure 9 shows the specific results. The trends exhibited by the data, as well as the relationships among the three methods, are largely consistent with those presented for DeepSeek-R1-Distill-Qwen-7B. Consequently, the final conclusions are also consistent.